it says here to rewrite each equation in standard form.
1. y=x+9 2. y=3x+7 3.y=4x+8
Please when answering let me know how to do it i really need help in this problem
Thanks

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

rewrite each equation in standard form.
Note: Standard form for a line is ax+by = c
-----
1. y=x+9
x-y = -9
--------------
2. y=3x+7
3x-y = -7
--------------
3.y=4x+8
4x-y=-8
===============
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi
standard form of a line is in the form Ax + By = C , where a is positive.
Performing the operations necessary to have this format:
1. y = x+9 |subtracting y and 9 from both sides of the equation
-9 = x - y |in the form Ax + By = C
OR x - y = -9
2. y = 3x + 7 |subtracting y and 7 from both sides of the equation
-7 = 3x - y
OR 3x - y = -7 |in the form Ax + By = C
3. y = 4x + 8 |subtracting y and 8 from both sides of the equation
-8 = 4x - y
OR 4x - y = -8 |in the form Ax + By = C